OpenStudy (anonymous):

An equation is shown below: 9x - y = -2 Part A: Explain how you will show all of the solutions that satisfy this equation. Part B: Determine three different solutions for this equation. Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent.

OpenStudy (precal):

Part A: solve for y Part B: Just pick any three points, I would use x=0, x=1, x=2 Part C: they are asking for parallel lines because inconsistent means no solution so write an equation that has the same slope

OpenStudy (nirmalnema):

9x - y = -2 y=9x+2 ... 2) x=1 y =11 x=2 y=20 x=3 y = 29 3)y=9x+c c = any no..except 2

OpenStudy (anonymous):

Thank youuu

OpenStudy (nirmalnema):

welcome.. :)